OpenStudy (sloppycanada):

The letters x and y represent rectangular coordinates. Write the following equation using polar coordinates (r, θ). x^2 + 4y^2 = 4

OpenStudy (sloppycanada):

Is my answer: \[\cos^2\theta+4\sin^2\theta=4r\]

OpenStudy (mathmale):

To engineer the necessary conversion: 1) replace "x" with "r cos theta." 2) replace "y" with "r sin theta." 3) Leave the constant (4) alone. Why multiply it by "r?"

OpenStudy (sloppycanada):

@mathmale Ohhhh okay. So it should be - \[r^2(\cos^2\theta +4\sin^2\theta) = 4\]